Parth (parthkohli):

@Limitless can you tell me what that capital pi meant? \(n!=\prod_{1 \leq f \leq n}f=1\cdot 2\cdot 3\cdots n\) Can I have more examples please?

OpenStudy (anonymous):

multiply

OpenStudy (anonymous):

just like \(\sum\) means add

Parth (parthkohli):

I don't get how that represents factorial. what's up with the f?

OpenStudy (anonymous):

\(f\) is the index of the productand. You multiply through all \(f\) in the set \(\{1,\dots,n\}\).

OpenStudy (anonymous):

the \(1\leq f\leq n\) tells you to multiply all the numbers together between 1 and n

Parth (parthkohli):

Hmmm I think I'm getting it.... more examples please?

OpenStudy (anonymous):

just like you could write \[\sum_{1\leq k\leq n}k\] to mean add all the numbers between one and n

OpenStudy (anonymous):

\[\prod_{1 \leq q \leq n}(q+1)=(1+1)(2+1)(3+1)\cdots(n+1)=(q+1)!\]

Parth (parthkohli):

\(k\) is constant, if I'm not wrong

OpenStudy (anonymous):

No, \(k\) is being used as an index.

OpenStudy (anonymous):

try here http://www.saddleback.edu/faculty/aorrison/mathhelp/sumprod1.htm

Parth (parthkohli):

What if I want to multiply all numbers between 1 and 10?

OpenStudy (anonymous):

Let \(n=10\) :P

Parth (parthkohli):

\( \color{Black}{\Rightarrow \prod_{1 \le f\le 10} f }\) Something like this?

OpenStudy (anonymous):

i.e. \[\prod_{1 \leq q \leq 10}q=1\cdot 2\cdot 3\cdot 4\cdot 5\cdot 6\cdot 7\cdot 8\cdot 9\cdot 10=. . .\]

OpenStudy (anonymous):

Yes. I lagged.

OpenStudy (anonymous):

yeah, but you can also write it as you do with a summation lower limit at the bottom, upper on top

OpenStudy (zarkon):

i prefer \[n!=\prod_{\begin{matrix} 1 \leq f \leq n \\f\in\mathbb{Z} \end{matrix} }f=1\cdot 2\cdot 3\cdots n\]

OpenStudy (anonymous):

\[ \prod_{n=1}^5\frac{n}{n-1}\]

Parth (parthkohli):

Why do we have that \(\le\)?

Parth (parthkohli):

Oh ok that lower and upper thing is a lot more easier.

OpenStudy (anonymous):

Zarkon, that's a weebit pedantic. \(f \in \mathbb{Z}\) is implied.

Parth (parthkohli):

That's how I used to do in summation.

OpenStudy (anonymous):

yeah in some circumstances

OpenStudy (zarkon):

not always

OpenStudy (anonymous):

exactly, it is the same only with products

OpenStudy (anonymous):

Not always, no. But in most cases, yes. Particularly elementary ones such as factorials. But, I have to say, there is a weebit of elegance to your definition.

OpenStudy (anonymous):

\[n!=\prod_{k=1}^nk\]

Parth (parthkohli):

I think I get it...... \( \color{Black}{\Rightarrow \Large \prod_{n = 1} ^{10} k}\) Is this to add all numbers between one and ten?

Parth (parthkohli):

Or maybe you'll put a \(n \epsilon \mathbb{Z} \) too.

OpenStudy (anonymous):

Parth, I am using what's basically a notation employed by computer scientist Donald E. Knuth. It is a much superior notation to the delimited form (what Satellite is showing you). It has much easier manipulations. It's a weebit confusing, at first, however. But the two notations are equivalent. The one I'm using, though, is more easily manipulated.

Parth (parthkohli):

I'm in trouble because I'll have to choose someone to give a medal to :P

OpenStudy (anonymous):

If you pick up a copy of Concrete Mathematics, he has an entire chapter dedicated to this. It's an elegant notation.

OpenStudy (anonymous):

i have enough, thank you. if i get any more they will take my bike away

OpenStudy (anonymous):

ima help u @ParthKohli :u give someone i give the other ;)

Parth (parthkohli):

Hmm so maybe Limitless is fine. Would you please not mind showing me how to write the product of numbers 1 to 10?

OpenStudy (anonymous):

I have already.

Parth (parthkohli):

Haha I want sat's notation.

OpenStudy (anonymous):

limitless is certainly correct about usage and indexing. later you will see infinite products, such as \[\prod_{p\in P} \frac{1}{1-p^{-s}} = \prod_{p\in P} \sum_{n=0}^{\infty} p^{-ns} = \sum_{n=1}^{\infty} \frac{1}{n^s} = \zeta(s) \]

OpenStudy (anonymous):

\[\prod_{k=1}^{10}k\]

Parth (parthkohli):

@Limitless where will I write k \(\epsilon \) \(\mathbb{Z}\)?

OpenStudy (anonymous):

you might wan to index over numbers other than integers. the one i wrote above is the riemann zeta function, and the index is over prime numbers

Parth (parthkohli):

I'll see the page that you linked me to. Thank you @satellite73 @Limitless !!!

OpenStudy (anonymous):

Parth, you're getting into another question now.

OpenStudy (zarkon):

you dont need \(k\in \mathbb{Z}\) with this notation \[\prod_{k=1}^{10}k\]

Parth (parthkohli):

Also @Zarkon